OpenStudy (larsa):

If cos θ =-5/13 and sin θ < 0, what is tan θ?

OpenStudy (bobo-i-bo):

You can find sin(theta) by using the identity \[(\cos(\theta))^{2}+(\sin(\theta))^{2}=1\]. Then you can find \[\tan(\theta)\] by the identity \[\tan(\theta)=\frac{ \sin(\theta) }{\cos(\theta) }\]

OpenStudy (anthonyym):

Well you can find theta using inverse cosine and a calculator. arccos(-5/13) = theta theta = 112.6198649 sin(112.6198649) = 0.923, that's less than 1. Then tan(112.6198649) = -2.4

OpenStudy (jdoe0001):

well.... hmmm so recall your SOH CAH TOA \(sin(\theta)=\cfrac{opposite}{hypotenuse} \qquad \qquad % cosine {\color{brown}{ cos(\theta)=\cfrac{adjacent}{hypotenuse} }} \\ \quad \\ % tangent tan(\theta)=\cfrac{opposite}{adjacent}\) so, notice the cosine identity, that means \(cos(\theta)=-\cfrac{5}{13}\implies \cfrac{adjacent}{hypotenuse}\) now, recall that, the hypotenuse is never negative, since it's just the radius so on -5/13 it can't be 5/-13 so, that means the -5 is really the adjacent side thus \(cos(\theta)=-\cfrac{5}{13}\implies \cfrac{adjacent}{hypotenuse}\implies \cfrac{adjacent=-5}{hypotenuse=13}\)